Subtract: (a – b) from (a + b) Subtract: (a – b) from (a + b)

Subtract:
 (a – b) from (a + b)
Subtract:
 (a – b) from (a + b)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

When term have the same algebraic factors, they are like terms.
Then, we have to subtract the like terms
= (a + b)–(a - b)
= a + b–a+ b
= a–a + b +b
= a (1 - 1) + b (1 + 1)
= a (0) + b (2)
= 2b

​Thanks